please help me solve this equation: +Translate+nine+less+then+five+times+a+number+is+16+into+an+equation+.
Answer by checkley77(12844)

5N-9=16
5N=16+9
5N=25
N=25/5
N=5
PROOF:
5*5-9=16
25-9=16
16=16